OpenStudy (anonymous):

How to find the value of r if 2pi(r)^2 + 14pi*r -231=0 This looks like a quadratic equation so I made x=pi*r My answer for r is 2.5 but in my textbook it has it as 3.5. Could someone please help? Thanks

OpenStudy (anonymous):

I get 2.5 also

OpenStudy (ranga):

I get x = 3.5 or x = -10.5 Divide the whole equation by 2pi r^2 + 7r -231/(2pi) = 0 or r^2 + 7r - 36.78 = 0 Discriminant = b^2 - 4ac = 7^2 + (4)(36.78) = 196. Square root = +/- 14 r = (-7 + 14) / 2 = 3.5 or r = (-7 - 14) / 3.5 = -10.5